# JWST Observations of the Double Nucleus in NGC 4486B: Possible Evidence for a Recent Binary SMBH Merger and Recoil

### v1

@@ 2.2 JWST NIRSpec-IFU kinematics
The kinematic datacube was Voronoi-tesselated (Cappellari.2003_voronoi) to achieve a uniform signal/noise ratio of $\sim 80$ across the NIRSpec field-of-view. Line-of-sight stellar kinematics were extracted using the penalized pixel-fitting method (pPXF; Cappellari.2023), fitting each Voronoi-binned spectrum with synthetic templates from the PHOENIX library (Husser.2013). We model the line-of-sight velocity distribution (LOSVD) using a Gauss-Hermite (GH) expansion as implemented in pPXF to extract mean velocity $v$, the width $\sigma$ of the best-fit Gaussian, and deviations captured by $h_{3}$ (related to skewness) and $h_{4}$ (related to kurtosis) (e.g., Marel.1993; Gerhard.1993).
@@
The left and middle panels of the first row in Figure 2 show $v_{\mathrm{los}}$ (in the galaxy rest-frame) and the velocity dispersion maps, respectively, obtained from JWST-NIRSpec. The white rectangle encloses the same spatial region in all three panels. The left-hand panel shows that the galaxy’s central velocity (vertical dashed line) lies between P2 (black dot) and P1 (purple dot), but is closer to P1. It is clear from the middle panel that the peak of the velocity dispersion (reddest bin) coincides with the fainter peak (black dot), which we therefore assume to be the position of the SMBH. The $\sigma$ peak, measured from NIRSpec, has a maximum value of $\sigma=287.2\pm 7.4\,\mathrm{km\ s}^{-1}$. This offset in the velocity dispersion peak is also seen in the second row of Figure 2, which shows 1D-kinematic profiles through the major axis of NGC 4486B. In this figure, the left side of the kinematic maps is mirror-reflected about the rotation axis to enable easier comparison of the two sides, and $|v|$ and $\sigma$ along the major axis are plotted. The vertical grey band shows the region encompassed by the white rectangle in Figure 2. In addition to the fact that the velocity dispersion profile is not symmetric about the center, we see in the top panel that $|v|$ of stars in the vicinity of the presumed SMBH is higher than for stars on the opposite side of the galaxy by $16\,\mathrm{km\ s}^{-1}$ The lowest panel shows the ratio $|v|/\sigma$, indicating that this galaxy has roughly equal amounts of rotational and dispersion support over the range of radii examined.
2.3 Schwarzschild model and $N$-body model for NGC 4486B
@@
To constrain the large-scale mass distribution, in Paper 1 we supplemented the JWST data with long-slit kinematics from CFHT/SIS (Kormendy.1997), using only the outer regions ($r\gtrsim 0.9\arcsec$) to minimize resolution-related discrepancies. The two-dimensional kinematic maps and the unsharp-masked image in Figure 2 reveal a well-resolved double nucleus and a velocity-dispersion peak offset from the photometric center. In Paper 1, we combined these long-slit kinematics with the photometric data described in Section 2.1 and the JWST kinematics (Section 2.2) to construct dynamical models of the galaxy and determine the SMBH mass. We applied two independent dynamical approaches: axisymmetric Schwarzschild orbit-superposition modeling (e.g. Schwarzschild.1979; vanderMarel.1998; Cretton.1999; Gebhardt.2003; Valluri.2004; Thomas.2004; vandenBosch.2010; Vasiliev.2020A) (which fits the full LOSVD including higher-order moments), and axisymmetric Jeans Anisotropic Modeling (JAM, Cappellari.2008) (which used only the first two velocity moments). We constrained the black hole mass as well as other properties of NGC 4486B. The two methods gave us broadly similar results and we refer readers to Paper 1 for further details. We restrict our discussion in this work to the orbit superposition models that were constructed with the FORSTAND code (Vasiliev.2020A).
@@
As both JAM and Schwarzschild modeling assume global symmetry and steady-state equilibrium, they cannot reproduce the pronounced asymmetries in the JWST-NIRSpec IFU kinematics and light distribution associated with the double nucleus. To gauge the impact of this modeling limitation, in Paper 1 we analyze three data configurations: (i) the original asymmetric kinematic maps; (ii) maps in which spaxels at and around the peak of the velocity dispersion, $\sigma$, are masked across all moments, suppressing the strongest asymmetry allowing us to probe the minimum black hole mass allowed; and (iii) a final test in which only the $\sigma$ map was modified by shifting its peak to the photometric center, an intentionally unphysical but informative symmetrization used to explore the maximum mass permitted under the symmetry assumption. These models yielded $2.8^{+0.6}_{-0.4}\times 10^{8}\leq{\cal M}_{BH}/\mbox{$M_{\odot}$}\leq 5.3^{+0.9}_{-0.9}\times 10^{8}$ with the detection significant at $3\sigma$ in all models. The inclusion or exclusion of a dark matter halo has a negligible impact on the measured ${\cal M}_{BH}$. Our Schwarzschild models, fit directly to the original JWST kinematics (no shifting or masking of the peak), yielded a black hole mass ${\cal M}_{BH}=3.6\pm{0.7}\times 10^{8}$$M_{\odot}$, a stellar mass-to-light ratio of $M/L_{[F850LP]}=3.9\pm 0.5$ corresponding to a stellar mass $M_{*}=7.0\pm 0.8\times 10^{9}\mbox{$M_{\odot}$}$, and an inclination $\theta\sim 70^{\circ}$.
@@
In addition to providing the globally optimal model parameters, the Schwarzschild method delivers the full orbital distribution function through the weights assigned to each orbit in the library. These weights allow us to generate a self-consistent, data-driven $N$-body realization of the galaxy. Although one of our best-fitting models includes a modest dark matter fraction ($\sim 14\%$ by mass), for the purposes of this study, we construct the $N$-body model using only the stellar component in order to simplify the subsequent dynamical experiments. The initial conditions for our simulations are drawn from the Schwarzschild model with ${\cal M}_{BH}=3.51\times 10^{8}\mbox{$M_{\odot}$}$, and total stellar mass $M_{*}=7.44\times 10^{9}\mbox{$M_{\odot}$}$, and an intrinsic axisymmetric shape with axis ratio $0.83$.
3 Evidence of an Eccentric Disk formed around a recoiling SMBH
@@
3.1 Comparison of nuclear structure and kinematics of NGC 4486B with eccentric nuclear disk simulations
@@
Wernke.2021 used $N$-body simulations to produce mock photometric and two-dimensional kinematic maps of ENDs for a wide range of viewing orientations. Their results show that the observed morphology depends strongly on the angle between the eccentricity vector (the disk’s major axis) and the line of sight. When the disk is viewed edge-on with its major axis aligned along the line of sight, only a single, centered nucleus is seen. If the disk is viewed edge-on with its minor axis along the line of sight, the nucleus appears offset. For edge-on or nearly edge-on views at intermediate angles relative to the eccentricity vector, the system appears as a double nucleus, where the fainter secondary peak lies close to the location of the black hole (see Figure 2 in Wernke.2021). They also found that the velocity dispersion is generally elevated and peaks at the location of the SMBH, irrespective of the viewing angle, and this dispersion peak typically does not coincide with the photometric center (brightest peak) unless the END is viewed such that it appears as a single centered nucleus.
@@
We therefore expect a double nucleus in NGC 4486B, consistent with our nearly edge-on viewing angle ($\theta\sim 70^{\circ}$), inferred from our Schwarzschild models. The morphology in the right-hand panel of Figure 2 closely resembles the mock ENDs viewed nearly edge-on at intermediate angles to the eccentricity vector in Wernke.2021, supporting this interpretation and consistent with the earlier conclusion of Lauer.1996. In such nearly edge-on ($\theta\sim 70^{\circ}$) configurations, the velocity-dispersion peak is expected to lie near the SMBH rather than at the brightest photometric peak—a behavior clearly seen in Figure 2 (first row, middle). A further examination of the 1D kinematic profiles of NGC 4486B along the major axis intersecting both nuclei (the second row of Figure 2) shows a clear asymmetry between the left and right sides of the galaxy: $|v|$ is elevated by $\sim 16\,\mathrm{km\ s}^{-1}$ and $\sigma$ is elevated by $\sim 57\,\mathrm{km\ s}^{-1}$ compared to the opposite side of the galaxy. Similar kinematic asymmetries and the offset of the velocity dispersion peak from the brighter peak strongly support the eccentric disk simulations (Wernke.2021), which also suggests local dynamical disturbances near the velocity dispersion peak (the presumed location of the SMBH). The enhanced $|v|$ may reflect streaming motions within the eccentric disk or, alternatively, residual motion of the SMBH relative to the center of the galaxy. Since the mock kinematics and images in Wernke.2021 were produced from simulations of isolated ENDs surrounding SMBH, whereas the one in NGC 4486B is embedded in an elliptical galaxy, additional simulations are needed to fully understand the offsets.
@@
As mentioned in Section 1, one of the proposed formation pathways for ENDs is a GW recoil of an SMBH following coalescence (Akiba.2021; Bright.2024). Such a recoil can drive stars that were initially on nearly circular prograde orbits onto eccentric trajectories and flip a substantial fraction, particularly at larger radii, into retrograde motion, as the displaced SMBH effectively outruns the outer stars and reverses their angular momenta in its rest frame.
@@
The resulting disk therefore contains a mixture of prograde and retrograde orbits, with the retrograde fraction increasing with radius; more powerful kicks generate a correspondingly larger retrograde population (Bright.2024). Simulations further indicate that this transformation occurs on very short timescales - within only a few dynamical times, $\sim 10^{4}-10^{5}$ yr on parsec scales (Akiba.2021). In cases where the disk hosts a substantial retrograde component, apsidal precession slows dramatically, allowing the lopsided eccentric disk structure to survive for $\gtrsim 10^{8}-10^{9}$ yr as a long-lived dynamical fossil of the recoil (Madigan.2018).
@@
Although our Schwarzschild dynamical models from Paper 1 do not explicitly model the double nucleus, the fact that they were designed to fit the observed kinematics provides a window into the orbital architecture of the nuclear region. We now examine the best-fit Schwarzschild model from Paper 1 to assess how the fraction of retrograde orbits varies with radius. For each orbit we extract mean radius $R$, and orbital circularity $\lambda=L_{z}/L_{\rm circ}$ (retrograde if $\lambda<0$) ($L_{z}$ is the angular momentum of the orbit and $L_{\rm circ}$ the angular momentum of a circular orbit with the same binding energy. We keep the orbits with $0\leq R\leq 1^{\prime\prime}$, then divide this range into 20 uniform bins. Since the orbit weights ($w$) in the Schwarzschild model correspond to the mass of stars on a given orbit, in each bin we compute the weighted share of retrograde orbits (retrograde weight divided by total weight in that bin). We compute uncertainties that reflect both the retrograde fraction and the number of effectively independent orbits contributing to a bin, reporting central 68% credible intervals using a Jeffreys prior binomial model (Jeffreys.1946) with an effective sample size derived from the orbit weights. To test robustness against discreteness noise and finite library size, we repeated the modeling in Paper 1 and measurement of the retrograde orbit fraction with an orbit library five times larger (100,000 orbits) than our standard model in that paper. The resulting profiles from the original and large-library model agree within the quoted uncertainties, indicating that our conclusions are not driven by sampling effects. We only show the results for the large-library model. The top panel of Figure 3 shows the retrograde-orbit fraction as a function of radius for the Schwarzschild model, while the bottom panel shows the corresponding orbital circularity distribution. The vertical dashed line at $R=0.25^{\prime\prime}$ ($\sim 20\,\mathrm{pc}$) marks the edge of the eccentric disk region in the white rectangle in Figure 2. Our results closely resemble the orbital distribution of an eccentric nuclear disk formed via a recoil kick (Akiba.2023), which also exhibits a substantial retrograde population ($\sim$50%), particularly in the outer regions of the disk. In Figure 14 of Akiba.2023, the top panel shows the orbital circularity distribution and the bottom panel shows the retrograde-orbit fraction. Although our Schwarzschild models assume axisymmetry and therefore cannot reproduce the full asymmetric structure of the eccentric disk, the emergence of a substantial retrograde component is an inevitable outcome of reproducing the observed JWST kinematics.
@@
Cretton.1999 demonstrated that Schwarzschild models can sometimes assign artificially large weights to retrograde orbits when the kinematics are constrained using only the GH expansion of the LOSVD. In their study of a rapidly rotating S0 galaxy, models without a central black hole reproduced the measured $h_{3}$ and $h_{4}$ moments by populating a large fraction of retrograde (counter-rotating) orbits, even though the observed velocity profiles showed no such component. This “spurious counter-rotation” arises because truncated GH series cannot uniquely describe the full LOSVD, allowing the model to exploit retrograde orbits as a fitting freedom. Importantly, this effect is most pronounced in cold, rotationally supported systems with $|v|/\sigma\gtrsim 1.5$, where small distortions in the LOSVD skewness can be mimicked by a retrograde tail.
@@
By contrast, NGC 4486B is a much more dispersion-dominated system: as shown in the bottom right panel of Figure 2, we measure $|v|/\sigma\lesssim 0.45$ in the inner regions, increasing only to $\sim 0.6$ at the outer edge of our JWST-NIRSpec field of view. In this regime, the degeneracies highlighted by Cretton.1999 are much less severe, making it unlikely that the strong retrograde signal we find is purely an artifact of the GH parameterization. Instead, the retrograde component appears to be an intrinsic dynamical requirement of the system.
@@
To test the importance of retrograde orbits explicitly, we re-ran our best-fit Schwarzschild models with identical initial conditions but excluded all retrograde orbits from the library. This increased the $\chi^{2}$ of the best-fit model by $\sim 23\%$, and the $\sigma$ map becomes vertically elongated and flattened along the projected minor axis. In contrast, the model including retrograde orbits produces a more rounded $\sigma$ structure, in much closer agreement with the observed data. Figure A1 in the Appendix compares the resulting two-dimensional $v$ and $\sigma$ maps: the first row shows the best-fit model with retrograde orbits, while the second row shows the best-fit model with all retrograde orbits removed. This confirms that the retrograde component is required to reproduce the kinematics of NGC 4486B. The presence of a significant retrograde orbit fraction in the region that appears to be dominated by the END thus offers tantalizing evidence that the disk formed as a result of a past SMBH recoil event.
@@
In addition to the recoiling black hole scenario, simulations show that the hardening of an SMBH binary can itself produce a large retrograde fraction in the nuclear orbital distribution: three-body slingshot interactions preferentially deplete low-$L$ and co-rotating (prograde) stars, leaving a tangentially biased remnant in which stable, counter-rotating orbits dominate inside the binary’s sphere of influence (e.g., Meiron.2010; Rantala.2019). Given that NGC 4486B hosts both a flat stellar core and a central SMBH, a binary-scouring scenario provides an alternative explanation for the large retrograde fraction we infer from the Schwarzschild model, offering complementary support for the possibility that the black hole in NGC 4486B is the result of the merger of a binary black hole.
3.2 Constraints on the Recoil Kick and SMBH Binary Parameters

### v2

@@ II.2 JWST NIRSpec-IFU kinematics
The kinematic datacube was Voronoi-tessellated (Cappellari & Copin, 2003) to achieve a uniform signal/noise ratio of $\sim 80$ across the NIRSpec field-of-view. Line-of-sight stellar kinematics were extracted using the penalized pixel-fitting method (pPXF; Cappellari, 2023), fitting each Voronoi-binned spectrum with synthetic templates from the PHOENIX library (Husser et al., 2013). We model the line-of-sight velocity distribution (LOSVD) using a Gauss-Hermite (GH) expansion as implemented in pPXF to extract mean velocity $v$, the width $\sigma$ of the best-fit Gaussian, and deviations captured by $h_{3}$ (related to skewness) and $h_{4}$ (related to kurtosis) (e.g., van der Marel & Franx, 1993; Gerhard, 1993).
@@
The left and middle panels of the first row in Figure 2 show $v_{\mathrm{los}}$ (in the galaxy rest-frame) and the velocity dispersion maps, respectively, obtained from JWST-NIRSpec. The white rectangle encloses the same spatial region in all three panels. The left-hand panel shows that the galaxy’s central velocity (vertical dashed line) lies between P2 (black dot) and P1 (purple dot), but is closer to P1. It is clear from the middle panel that the peak of the velocity dispersion (reddest bin) coincides with the fainter peak (black dot), which we therefore assume to be the position of the SMBH. The $\sigma$ peak, measured from NIRSpec, has a maximum value of $\sigma=287.2\pm 7.4\,\mathrm{km\ s}^{-1}$. This offset in the velocity dispersion peak is also seen in the second row of Figure 2, which shows 1D-kinematic profiles through the major axis of NGC 4486B. In this figure, the left side of the kinematic maps is mirror-reflected about the rotation axis to enable easier comparison of the two sides, and $|v|$ and $\sigma$ along the major axis are plotted. The vertical grey band shows the region encompassed by the white rectangle in Figure 2. In addition to the fact that the velocity dispersion profile is not symmetric about the center, we see in the top panel that $|v|$ of stars in the vicinity of the presumed SMBH is higher than for stars on the opposite side of the galaxy by $16\,\mathrm{km\ s}^{-1}$ The rightmost panel shows the ratio $|v|/\sigma$, indicating that this galaxy has roughly equal amounts of rotational and dispersion support over the range of radii examined.
Taken together, the resolved double nucleus, the offset of the velocity-dispersion peak from the photometric maximum, and the asymmetric major-axis $|v|$ and $\sigma$ profiles shown in Figure 2 provide purely observational photometric and kinematic signatures that are most naturally explained by an END in NGC 4486B.
II.3 Schwarzschild model and $N$-body model for NGC 4486B
In Paper 1, we combined the photometric data described in Section II.1, the JWST kinematics summarized in Section II.2, and supplementary long-slit CFHT/SIS measurements from Kormendy et al. (1997) to construct dynamical models of NGC 4486B and constrain both the black hole mass and the large-scale mass distribution.
@@ II.3 Schwarzschild model and $N$-body model for NGC 4486B
The two-dimensional JWST kinematic maps and the unsharp-masked image shown in Figure 2—presented here for completeness—reveal a well-resolved double nucleus and a velocity-dispersion peak offset from the photometric center. In Paper 1, we applied two independent dynamical approaches to these data: axisymmetric Schwarzschild orbit-superposition modeling (e.g. Schwarzschild, 1979; van der Marel et al., 1998; Cretton et al., 1999; Gebhardt et al., 2003; Valluri et al., 2004; Thomas et al., 2004; van den Bosch & de Zeeuw, 2010; Vasiliev & Valluri, 2020), which fits the full LOSVD including higher-order moments, and axisymmetric Jeans Anisotropic Modeling (JAM; Cappellari, 2008), which uses only the first two velocity moments. Both approaches yielded broadly consistent SMBH mass estimates. We refer the reader to Paper 1 for full details of these models and results.
In the present work, we restrict our discussion to the orbit-superposition models constructed with the FORSTAND code (Vasiliev & Valluri, 2020), using them to examine the orbital structure and implications for the formation of the eccentric nuclear disk.
@@
As both JAM and Schwarzschild modeling assume global symmetry and steady-state equilibrium, they cannot reproduce the pronounced asymmetries in the JWST-NIRSpec IFU kinematics and light distribution associated with the double nucleus. To gauge the impact of this modeling limitation, in Paper 1 we analyze three data configurations: (i) the original asymmetric kinematic maps; (ii) maps in which spaxels at and around the peak of the velocity dispersion, $\sigma$, are masked across all moments, suppressing the strongest asymmetry allowing us to probe the minimum black hole mass allowed; and (iii) a final test in which only the $\sigma$ map was modified by shifting its peak to the photometric center, an intentionally unphysical but informative symmetrization used to explore the maximum mass permitted under the symmetry assumption. These models yielded $2.8^{+0.6}_{-0.4}\times 10^{8}\leq{\cal M}_{BH}/\mbox{$M_{\odot}$}\leq 5.3^{+0.9}_{-0.9}\times 10^{8}$, the detection significant at $3\sigma$ in all models. The inclusion or exclusion of a dark matter halo has a negligible impact on the measured ${\cal M}_{BH}$. Our Schwarzschild models, fit directly to the original JWST kinematics (no shifting or masking of the peak), yielded a black hole mass ${\cal M}_{BH}=3.6\pm{0.7}\times 10^{8}$$M_{\odot}$, a stellar mass-to-light ratio of $M/L_{[F850LP]}=3.9\pm 0.5$ corresponding to a stellar mass $M_{*}=7.0\pm 0.8\times 10^{9}\mbox{$M_{\odot}$}$, and an inclination $\theta\sim 70^{\circ}$.
In addition to providing the globally optimal model parameters, the Schwarzschild method delivers the full orbital distribution function through the weights assigned to each orbit in the library. While the black hole mass and stellar mass reported above represent marginalized estimates (median and uncertainties) over the full model ensemble, the construction of an $N$-body realization requires selecting a single best-fitting model corresponding to the minimum $\chi^{2}$. These weights allow us to generate a self-consistent, data-driven $N$-body realization of the galaxy.
@@
Although one of our best-fitting models includes a modest dark matter fraction ($\sim 14\%$ by mass), for the purposes of this study, we construct the $N$-body model using only the stellar component in order to simplify the subsequent dynamical experiments. The initial conditions for our simulations are drawn from the minimum-$\chi^{2}$ Schwarzschild model, with ${\cal M}_{BH}=3.51\times 10^{8}\mbox{$M_{\odot}$}$, mass-to-light ratio $\Upsilon_{[F850LP]}=4.1$, a total stellar mass $M_{*}=7.44\times 10^{9}\mbox{$M_{\odot}$}$, and an intrinsic axisymmetric shape with axis ratio $0.83$. These values differ slightly from the marginalized medians but lie well within the quoted uncertainties.
We emphasize that this $N$-body realization is based on the axisymmetric Schwarzschild model and therefore does not explicitly contain an END. The purpose of these simulations is solely to trace the large-scale trajectory and damping timescale of a recoiling SMBH in the global potential of NGC 4486B. They are not intended to model the formation or internal structure of the END. A fully self-consistent, non-axisymmetric $N$-body treatment tailored to the observed eccentric disk structure is beyond the scope of the present work and will be explored in future studies.
III Formation of an Eccentric Nuclear Disk from a Recoiling SMBH
@@
III.1 Comparison of nuclear structure and kinematics of NGC 4486B with eccentric nuclear disk simulations
@@
Wernke & Madigan (2021) used $N$-body simulations to produce mock photometric and two-dimensional kinematic maps of ENDs for a wide range of viewing orientations. Their results show that the observed morphology depends strongly on the angle between the eccentricity vector (the disk’s major axis) and the line of sight. When the disk is viewed edge-on with its major axis aligned along the line of sight, only a single, centered nucleus is seen. If the disk is viewed edge-on with its minor axis along the line of sight, the nucleus appears offset. For edge-on or nearly edge-on views at intermediate angles relative to the eccentricity vector, the system appears as a double nucleus, where the fainter secondary peak lies close to the location of the black hole (see Figure 2 in Wernke & Madigan, 2021). They also found that the velocity dispersion is generally elevated and peaks at the location of the SMBH, irrespective of the viewing angle, and this dispersion peak typically does not coincide with the photometric center (brightest peak) unless the END is viewed such that it appears as a single centered nucleus.
@@
We therefore expect a double nucleus in NGC 4486B, consistent with our nearly edge-on viewing angle ($\theta\sim 70^{\circ}$), inferred from our Schwarzschild models. The morphology in the right-hand panel of Figure 2 closely resembles the mock ENDs viewed nearly edge-on at intermediate angles to the eccentricity vector in Wernke & Madigan (2021), supporting this interpretation and consistent with the earlier conclusion of Lauer et al. (1996). In such nearly edge-on ($\theta\sim 70^{\circ}$) configurations, the velocity-dispersion peak is expected to lie near the SMBH rather than at the brightest photometric peak—a behavior clearly seen in Figure 2 (first row, middle). A further examination of the 1D kinematic profiles of NGC 4486B along the major axis intersecting both nuclei (the second row of Figure 2) shows a clear asymmetry between the left and right sides of the galaxy: $|v|$ is elevated by $\sim 16\,\mathrm{km\ s}^{-1}$ and $\sigma$ is elevated by $\sim 57\,\mathrm{km\ s}^{-1}$ compared to the opposite side of the galaxy. Similar kinematic asymmetries and the offset of the velocity dispersion peak from the brighter peak strongly support the eccentric disk simulations (Wernke & Madigan, 2021), which also suggests local dynamical disturbances near the velocity dispersion peak (the presumed location of the SMBH). The enhanced $|v|$ may reflect streaming motions within the eccentric disk or, alternatively, residual motion of the SMBH relative to the center of the galaxy. Since the mock kinematics and images in Wernke & Madigan (2021) were produced from simulations of isolated ENDs surrounding SMBH, whereas the one in NGC 4486B is embedded in an elliptical galaxy, additional simulations are needed to fully understand the offsets.
@@
As mentioned in Section I, one of the proposed formation pathways for ENDs is a GW recoil of an SMBH following coalescence (Akiba & Madigan, 2021; Bright et al., 2024). Such a recoil can drive stars that were initially on nearly circular prograde orbits onto eccentric trajectories and flip a substantial fraction, particularly at larger radii, into retrograde motion, as the displaced SMBH effectively outruns the outer stars and reverses their angular momenta in its rest frame.
@@
The resulting disk therefore contains a mixture of prograde and retrograde orbits, with the retrograde fraction increasing with radius; more powerful kicks generate a correspondingly larger retrograde population (Bright et al., 2024). Simulations further indicate that this transformation occurs on very short timescales - within only a few dynamical times, $\sim 10^{4}-10^{5}$ yr on parsec scales (Akiba & Madigan, 2021). In cases where the disk hosts a substantial retrograde component, apsidal precession slows dramatically, allowing the lopsided eccentric disk structure to survive for $\gtrsim 10^{8}-10^{9}$ yr as a long-lived dynamical fossil of the recoil (Madigan et al., 2018).
Before proceeding, we emphasize that the analysis presented below is not intended to establish the existence of the END itself. The observational evidence for an END in NGC 4486B is provided directly by the photometric and kinematic signatures discussed in Section II and shown in Figure 2. Instead, the purpose of the Schwarzschild-based analysis presented here is to examine the orbital phase-space structure required by the observed kinematics and to assess whether it is consistent with the GW recoil kick-based END formation scenarios.
@@
Although our Schwarzschild dynamical models from Paper 1 do not explicitly model the double nucleus, the fact that they were designed to fit the observed kinematics provides a window into the orbital architecture of the nuclear region. We now examine the best-fit Schwarzschild model from Paper 1 to assess how the fraction of retrograde orbits varies with radius. For each orbit we extract mean radius $R$, and orbital circularity $\lambda=L_{z}/L_{\rm circ}$ (retrograde if $\lambda<0$), $L_{z}$ is the angular momentum of the orbit and $L_{\rm circ}$ the angular momentum of a circular orbit with the same binding energy. We keep the orbits with $0\leq R\leq 1^{\prime\prime}$, then divide this range into 20 uniform bins. Since the orbit weights ($w$) in the Schwarzschild model correspond to the mass of stars on a given orbit, in each bin we compute the weighted share of retrograde orbits (retrograde weight divided by total weight in that bin). We compute uncertainties that reflect both the retrograde fraction and the number of effectively independent orbits contributing to a bin, reporting central 68% credible intervals using a Jeffreys prior binomial model (Jeffreys, 1946) with an effective sample size derived from the orbit weights. To test robustness against discreteness noise and finite library size, we repeated the modeling in Paper 1 and measurement of the retrograde orbit fraction with an orbit library five times larger (100,000 orbits) than our standard model in that paper. The resulting profiles from the original and large-library model agree within the quoted uncertainties, indicating that our conclusions are not driven by sampling effects. We only show the results for the large-library model. The top panel of Figure 3 shows the retrograde-orbit fraction as a function of radius for the Schwarzschild model, while the bottom panel shows the corresponding orbital circularity distribution. The vertical dashed line at $R=0.25^{\prime\prime}$ ($\sim 20\,\mathrm{pc}$) marks the edge of the eccentric disk region in the white rectangle in Figure 2. Our results closely resemble the orbital distribution of an eccentric nuclear disk formed via a recoil kick (Akiba & Madigan, 2023), which also exhibits a substantial retrograde population ($\sim$50%), particularly in the outer regions of the disk. In Figure 14 of Akiba & Madigan (2023), the top panel shows the orbital circularity distribution and the bottom panel shows the retrograde-orbit fraction. Although our Schwarzschild models assume axisymmetry and therefore cannot reproduce the full asymmetric structure of an eccentric disk, they robustly demonstrate that the observed JWST kinematics require a substantial retrograde orbital component. This comparison is therefore used to assess consistency with proposed END formation scenarios, rather than to establish the existence of the END itself.
@@
Cretton et al. (1999) demonstrated that Schwarzschild models can sometimes assign artificially large weights to retrograde orbits when the kinematics are constrained using only the GH expansion of the LOSVD. In their study of a rapidly rotating S0 galaxy, models without a central black hole reproduced the measured $h_{3}$ and $h_{4}$ moments by populating a large fraction of retrograde (counter-rotating) orbits, even though the observed velocity profiles showed no such component. This “spurious counter-rotation” arises because truncated GH series cannot uniquely describe the full LOSVD, allowing the model to exploit retrograde orbits as a fitting freedom. Importantly, this effect is most pronounced in cold, rotationally supported systems with $|v|/\sigma\gtrsim 1.5$, where small distortions in the LOSVD skewness can be mimicked by a retrograde tail.
@@
By contrast, NGC 4486B is a much more dispersion-dominated system: as shown in the bottom right panel of Figure 2, we measure $|v|/\sigma\lesssim 0.45$ in the inner regions, increasing only to $\sim 0.6$ at the outer edge of our JWST-NIRSpec field of view. In this regime, the degeneracies highlighted by Cretton et al. (1999) are much less severe, making it unlikely that the strong retrograde signal we find is purely an artifact of the GH parameterization. Instead, the retrograde component appears to be an intrinsic dynamical requirement of the system.
@@
To test the importance of retrograde orbits explicitly, we re-ran our best-fit Schwarzschild models with identical initial conditions but excluded all retrograde orbits from the library. This increased the $\chi^{2}$ of the best-fit model by $\sim 23\%$, and the $\sigma$ map becomes vertically elongated and flattened along the projected minor axis. In contrast, the model including retrograde orbits produces a more rounded $\sigma$ structure, in much closer agreement with the observed data. Figure A1 in the Appendix compares the resulting two-dimensional $v$ and $\sigma$ maps: the first row shows the best-fit model with retrograde orbits, while the second row shows the best-fit model with all retrograde orbits removed. This confirms that the retrograde component is required to reproduce the kinematics of NGC 4486B. The presence of a significant retrograde orbit fraction in the region that appears to be dominated by the END, thus offers tantalizing evidence that the observed END is consistent with formation via a past SMBH recoil event.
@@
In addition to the recoiling black hole scenario, simulations show that the hardening of an SMBH binary can itself produce a large retrograde fraction in the nuclear orbital distribution: three-body slingshot interactions preferentially deplete low-$L$ and co-rotating (prograde) stars, leaving a tangentially biased remnant in which stable, counter-rotating orbits dominate inside the binary’s sphere of influence (e.g., Meiron & Laor, 2010; Rantala et al., 2019). Given that NGC 4486B hosts both a flat stellar core and a central SMBH, a binary-scouring scenario provides an alternative explanation for the large retrograde fraction we infer from the Schwarzschild model, offering complementary support for the possibility that the black hole in NGC 4486B is the result of the merger of a binary black hole.
III.2 Constraints on the Recoil Kick and SMBH Binary Parameters